OpenStudy (anonymous):

The value of the derivative of: y= 0.5x^3+2x^2+3x+1500 at x=5 is.....

OpenStudy (anonymous):

y' = 1.5x^2+4x+3 y'(3) = 1.5(3)^2+4(3)+3 y'(5) = 45+20+3 = 68 I think I did that math right. Please double check it because I did it in my head.

OpenStudy (anonymous):

put x=5 in 1.5x^2 + 4x + 3

OpenStudy (anonymous):

I reposted. I still have 3's in there. But the math is done right now. Sorry

OpenStudy (anonymous):

thanks for helping me..^^

OpenStudy (anonymous):

You're welcome. Do you understand how to differentiate?

OpenStudy (anonymous):

and the answer would be 60.5. @teic85 i did not understand your calculations :O 1.5*(5^2)= 45?

OpenStudy (anonymous):

do you believe 1.5*9 gives 45?

OpenStudy (anonymous):

lol sorry, i'm tired and im not using paper and pencil

OpenStudy (anonymous):

you did the whole answer wrong!!!! put x=5 in the equation!!

OpenStudy (anonymous):

Once you have y', you can just plug in the 5.

OpenStudy (anonymous):

I'm 99% my y' is correct.

OpenStudy (anonymous):

\[y' = 1.5x^2+4x+3 \]\[y'(5) = 1.5(5)^2+4(5)+3 \] \[y'(5) = 37.5+20+3 = 62.5 \]